solve this equation by sustitution
x=8-4y
2x-3y=13

Answer by mananth(16946)

2x-3y = 13 .............1

4y=8-x ---------2
/4
y=2-0.25x
Plug the value of y in (1)
2x-3(2-0.25x)=13
2x-6+0.75x=13
2x+0.75 x=13+6
2.75 x = 19
/ 2.75
x= 6.91
Plug the value of x in (1)
2x-3y=13
2*6.91-3y=13
13.82 -3y=13
-3y=-0.82
/-3
y=0.27